OpenStudy (anonymous):

Solve the following system using the substitution method. x + 6y = –5 2x – 9y = –17

OpenStudy (vishweshshrimali5):

multiply first eq. by 2 both sides 2x + 12y = -10 now 2x + 12y = -10 2x – 9y = –17 _______________ - + + 0 + 21y = 7 => y = 1/3 now put this value in any of the 2 eq. and solve for x

OpenStudy (anonymous):

let x=-(5+6Y)...(1) the substitute in the equation 2x -9y=-17...(2) where it will become 2(-5-6y) -9y=-17 the solve for y after finding y subtitute y to equation 1.